Suppose the demand for banana of AMU students is given by the demand function Qb=1200-9.5 pb + 16.2 p m + 0.2 Y, where Qb is the quantity of banana demanded ; pb is the price of banana in birr per unit ; pm is the price of mango in birr per unit and Y is the income of the student. Assume that pb is initially 45 birr per unit, Y is 100 birr and Qb is 127,500 per week, calculate
A) the price elasticity of demand and comment on type of elasticity
B) the income of demand and comment on type of the good
C) the cross-elasticity of demand and identify the type of the commodities
1
Expert's answer
2021-09-28T13:18:20-0400

A) Qb'=-9.5; E=-9.5×45/127,500=-0.003=-0.3% The demand is inelastic

B) Q'=0.2; E=0.2×100/127,500=0.00015=0.015%. The good is normal (E>0)

C) Q'=16.2; Pm can be found from the equation

127,500=1200-9.5×45+16.2Pm+0.2×100; Pm=7821.45

Then E=16.2×7821.45/127,500=0.9937=99.37% The goods are substitutes